OpenStudy (anonymous):

what is the quadratic equation for 100-121a^2

OpenStudy (anonymous):

The quadratic formula is: \[x = (-b +- \sqrt{b^2-4ac})/2a\] and the standard quadratic form is: ax^2 + bx + c = 0 So we can rewrite your equation as: -121(a^2) + 100(b) + c = 0 and for the quadratic formula we get: a = -121 b = 100 c = 0 Simply substitute and solve for X